If I purchased items for a fundraiser for $2.00 each and sold them for $4.00 each what is my profit percentage ?
Answer by jim_thompson5910(35256)

Difference: 4-2 = 2

Divide the difference by the initial cost: 2/2 = 1

Then multiply by 100: 1*100 = 100%

The profit percentage is 100%